# EEG-ImageNet: An Electroencephalogram Dataset and Benchmarks with Image Visual Stimuli of Multi-Granularity Labels

### v1

@@ Abstract
Identifying and reconstructing what we see from brain activity gives us a special insight into investigating how the biological visual system represents the world. While recent efforts have achieved high-performance image classification and high-quality image reconstruction from brain signals collected by Functional Magnetic Resonance Imaging (fMRI) or magnetoencephalogram (MEG), the expensiveness and bulkiness of these devices make relevant applications difficult to generalize to practical applications. On the other hand, Electroencephalography (EEG), despite its advantages of ease of use, cost-efficiency, high temporal resolution, and non-invasive nature, has not been fully explored in relevant studies due to the lack of comprehensive datasets. To address this gap, we introduce EEG-ImageNet, a novel EEG dataset comprising recordings from 16 subjects exposed to 4000 images selected from the ImageNet dataset. EEG-ImageNet consists of 5 times EEG-image pairs larger than existing similar EEG benchmarks. EEG-ImageNet is collected with image stimuli of multi-granularity labels, i.e., 40 images with coarse-grained labels and 40 with fine-grained labels. Based on it, we establish benchmarks for object classification and image reconstruction. Experiments with several commonly used models show that the best models can achieve object classification with accuracy around 60% and image reconstruction with two-way identification around 64%. These results demonstrate the dataset’s potential to advance EEG-based visual brain-computer interfaces, understand the visual perception of biological systems, and provide potential applications in improving machine visual models.
@@ 1 Introduction
Recent advancements in reconstructing visual experiences from the human brain have seen significant progress, largely driven by the extensive use of functional magnetic resonance imaging (fMRI) ([8, 22, 23]) and magnetoencephalogram (MEG) [3] datasets. fMRI and MEG are widely used to investigate various cognitive functions, neurological disorders, and brain connectivity patterns ([2, 40, 37, 35]). Driven by the use of deep neural networks, particularly diffusion-based and transformer-based models( [34, 30, 25, 6]), it is even possible to reconstruct human’s visual perceptions from fMRI or MEG recordings. The application of large-scale deep neural networks in neuroscience research has further underscored the importance of large-scale and high-quality datasets [27, 20]. For example, the Natural Scenes Dataset ([1]), contains up to hundreds of thousands of high-quality natural image-fMRI pairs of 8 subjects, providing a solid data foundation for recent work in visual neuroscience. These models and large-scale datasets, in turn, have opened new avenues for understanding the brain’s intricate functions and for developing advanced applications in brain-computer interfaces, neuroimaging, and beyond [33].
@@
On the other hand, electroencephalography (EEG) is another vital tool in neuroscience research. In comparison with fMRI and MEG, EEG is easy to use, cost-efficient, and has superior temporal resolution, making it a valuable tool for capturing rapid and real-time brain dynamics on the order of milliseconds [36]. EEG signals can be obtained non-invasively by placing electrodes on the scalp, making it a less intrusive method for monitoring brain activity. These attributes position EEG as another promising modality for visual neuroscience research.
Although visual reconstruction has been achieved using fMRI and MEG, the high cost and inconvenience of these devices limit their widespread application in practical settings. In contrast, EEG presents advantages over both fMRI and MEG with its cost-efficient and portable features. However, studies on visual perception with EEG signals are limited because of two challenges: (1) the lack of large-scale, high-quality EEG datasets and (2) existing EEG datasets typically featured coarse-grained image categories, lacking fine-grained categories. To the best of our knowledge, the most frequently used dataset is the data set provided by Spampinato et al. [32], which involves 6 participants each watching 2000 image stimuli. However, this dataset’s scale is smaller than existing fMRI datasets, limiting the possibilities for investigating neural aspects related to visual perception and developing deep learning models for relevant visual classification and reconstruction tasks.
On the other hand, the labels in existing EEG datasets are frequently coarse and lack the granularity needed for detailed analysis. Multi-granularity labels are essential because they allow for a more nuanced analysis at different levels of detail. For instance, labels can range from broad categories like “panda” or “golf ball” to more specific attributes like “Rottweiler” or “Samoyed”. These challenges underscore the necessity for new, large-scale EEG datasets with high-quality, multi-granularity labels. Such datasets would enable researchers to explore the intricacies of visual processing with greater accuracy and depth, facilitating advancements in both basic neuroscience and applied fields like brain-computer interfaces and clinical diagnostics.
To address these challenges, we present EEG-ImageNet, a novel EEG dataset specifically designed to promote research related to visual neuroscience, biomedical engineering, etc. EEG-ImageNet is a comprehensive dataset that includes EEG recordings from 16 subjects, each exposed to 4,000 images sourced from the ImageNet-21k [28]. These images span 80 different categories, with 50 images per category. The dataset is structured to support multi-granularity analysis, with 40 categories dedicated to coarse-grained tasks and 40 to fine-grained tasks. This dataset aims to bridge the gap in the current research landscape by providing a resource that leverages the strengths of EEG and supports diverse research needs. To demonstrate the utility of this dataset, we establish benchmarks for two primary tasks: object classification and image reconstruction. For the object classification task, we evaluated the dataset using several commonly used models, achieving a best accuracy of 60.88% on the 80-class classification. In the image reconstruction task, our experiments with advanced generative models produced promising results, with the best model achieving a two-way identification of 64.67%. These benchmarks demonstrate the dataset’s potential for advancing EEG-based research. Figure 1 shows the overall procedure of our dataset construction and benchmark design. By addressing the lack of large-scale, high-quality EEG datasets, EEG-ImageNet aims to drive EEG-based visual research forward, improve machine learning models, and provide deeper insights into visual perception and processing.
@@ 2 Related Work
In this section, we review some datasets related to visual recognition and neuroscience and compare them with EEG-ImageNet, as shown in Table 1.
2.1 Visual Recognition Dataset
Visual recognition is a cornerstone of computer vision, driven by datasets like ImageNet [28], CIFAR [16], and MS COCO [20]. ImageNet contains over 14 million annotated images, CIFAR-10/100 consist of thousands of 32x32 images in 10 and 100 classes, respectively, and MS COCO is known for its rich annotations supporting tasks such as object detection and segmentation. Later works have matched image datasets with additional modalities.
@@ 2.1 Visual Recognition Dataset
Efforts to combine visual recognition with neuroscience have led to datasets like the Microsoft COCO Captions [5], which adds textual descriptions to the MS COCO images, enhancing the dataset with a multimodal aspect for evaluating image captioning models. The SALICON dataset [10] extends MS COCO with eye-tracking data, enabling the study of visual attention and saliency through large-scale annotations. Neuroscience datasets utilizing fMRI have further enriched this field. The BOLD5000 dataset [4] includes fMRI data from subjects viewing 5000 images, aiding the exploration of visual perception. The Generic Object Decoding dataset [9] captures brain activity while subjects view and imagine objects, facilitating the decoding of mental images. The VIM-1 dataset [14] from the study “Identifying Natural Images from Human Brain Activity” demonstrates the feasibility of decoding viewed images from brain activity. Additionally, the NSD [1] is a large-scale fMRI dataset in visual neuroscience, recording high-resolution (1.8-mm) whole-brain 7T fMRI data from eight subjects exposed to 9,000–10,000 color natural scenes from the MS COCO dataset over the course of a year. Integrating neuroimaging and eye-tracking datasets with visual recognition tasks has opened new avenues for understanding how the brain interprets visual information, leading to insights and applications in brain-computer interfaces and neural decoding.
| Dataset | #Subjects | Modalities | Visual Stimuli | #Stimuli | #Stimuli per Subject |
| SALICON [10] | - | Eye-tracking | MS COCO [20] | 20,000 | - |
| BOLD5000 [4] | 4 | fMRI | ImageNet [28], MS COCO, SceneUN [38] | 5,254 | 1,157-1,798 |
| GOD [9] | 5 | fMRI | ImageNet | 1,200 | 1,200 |
| VIM-1 [14] | 2 | fMRI | Corel Stock Photo Library [11] | 1,870 | 1,870 |
| NSD [1] | 8 | fMRI, Eye-tracking | MS COCO | 9,000-10,000 | 22,000–30,000 |
| Spampinato et al. [32] | 6 | EEG | ImageNet | 2000 | 2000 |
| EEG-SVRec [42] | 30 | EEG, ECG | short videos | 2636 | 121.9 |
| EEG-ImageNet | 16 | EEG | ImageNet | 4,000 | 4,000 |
Note: The original SALICON dataset lacks information on the number of subjects involved.
2.2 EEG dataset
@@
fMRI is renowned for its high spatial resolution, allowing researchers to obtain detailed images of brain activity by measuring changes in blood flow [23]. In contrast, EEG offers several distinct advantages over fMRI. EEG is relatively easy to use and cost-efficient, with a straightforward setup that involves placing electrodes on the scalp to measure electrical activity. One of the most significant benefits of EEG is its exceptional temporal resolution, which captures neural dynamics on the order of milliseconds [36]. This high temporal resolution makes EEG ideal for studying fast-occurring brain processes and real-time neural responses, providing insights into the timing and sequence of neural events [21]. EEG’s non-invasive nature also makes it suitable for a wider range of participants. However, EEG signals collected using portable devices often have a low signal-to-noise ratio, which can complicate data analysis and reduce the accuracy of the results [13].
@@
Existing EEG datasets span a variety of research areas. The SEED [43] focuses on emotion recognition with detailed EEG recordings from subjects exposed to various emotional stimuli. The BCI Competition IV datasets [41] provide EEG data for motor imagery tasks, while the TUH EEG Corpus [31] is a large clinical EEG collection often used for benchmarking EEG data quality across different conditions. The DEAP [15] collects EEG and peripheral physiological signals from 32 participants as they watch 40 one-minute music videos, providing comprehensive emotional responses. Similarly, the AMIGOS [24] captures EEG and physiological responses from participants watching short video clips designed to evoke specific emotional states. In the realm of visual recognition, datasets like the EEG-Classification dataset [32] involve 6 subjects viewing 2,000 images across 40 object classes from the ImageNet10k. Another significant dataset is the EEG-SVRec [42], which includes EEG recordings from 30 participants interacting with short videos, aiming to capture detailed affective experiences.
In comparison, the EEG-ImageNet dataset offers several advantages. Firstly, its larger scale is conducive to training deep learning models. Secondly, the greater number of subjects facilitates inter-subject experiments and analyses. Additionally, the dataset features multi-granularity image labels, and the images are of high quality, selected from ImageNet21k and filtered to exclude small, blurry, or watermarked pictures.
@@ 3 Dataset Construction
During the data collection process of our user study, participants are presented with a visual stimuli dataset containing 4000 natural images from ImageNet. Throughout this process, we continuously record their EEG signals. The whole experimental process is carried out in the laboratory environment. This section describes the entire process of EEG-ImageNet dataset construction. The EEG-ImageNet dataset can be accessed openly through the url https://github.com/Promise-Z5Q2SQ/EEG-ImageNet-Dataset.
3.1 Ethical and Privacy
To protect participants’ privacy and physical health, our user study adheres to strict ethical guidelines for human research, with approval from the ethics committee of the School of Psychology at Tsinghua University. The study has undergone a comprehensive ethical review to safeguard participants’ rights. In accordance with ethical standards, we have taken several steps to protect participants’ privacy, including data anonymization and obtaining informed consent from all participants. Additionally, participants are thoroughly informed about the study’s objectives, procedures, and potential outcomes. The EEG data collection method employed in this research is non-invasive and poses no risk to participants. This approach ensures compliance with ethical standards while maintaining the integrity of the research findings.
@@ 3.2 Participants
We enlist a total of 16 participants via social media, including 10 males and 6 females. These participants are all college students aged between 21 and 27, with an average age of 24.06 and a standard deviation of 1.69. Their majors encompass computer science, mechanical engineering, chemistry, and environmental engineering, and they range from undergraduate to postgraduate levels. All participants are right-handed and assert their proficiency in utilizing image search engines in their daily routines. Each participant dedicates approximately 2 hours to complete the experiment, which includes 30 minutes for equipment setup and task instructions. Before the experiment, participants are informed of a compensation of US$11.8 per hour upon completion, to ensure the quality of the data collected for the study.
@@ 3.3 Stimuli Dataset
Among all categories, the first half is consistent with the EEG-Classification dataset ([32]), comprising 40 significantly distinct categories from ImageNet1k. We treat these as coarse-grained tasks. The latter 40 categories are designed as a fine-grained task, divided into 5 groups with 8 categories each. The categories within the same group share the same parent node in WordNet, and each category label is either a leaf node or a sub-leaf node in WordNet. This selection ensures that the chosen categories represent similar granularity while avoiding overly obscure categories, thereby minimizing potential biases in the experimental results. For instance, coarse-grained categories include items such as African elephants, pandas, mobile phones, golf balls, bananas, and pizzas. Under the parent node "musical instruments," the fine-grained categories include accordions, cellos, flutes, oboes, snare drums, and trombones. Detailed information about all the visual stimuli categories and their respective WordNet IDs can be found in our GitHub repository.
@@ 3.4 Procedure
Before engaging in the user study, participants are required to fill out an entry questionnaire and sign a consent about the protection of privacy security. They will receive an orientation regarding the primary tasks and operational procedures. Additionally, they will be notified of their right to withdraw from the study at any point. Before the main trials, participants will undergo a series of training trials designed to acquaint participants with the procedures of the formal experiments.
Every participant is required to select a random seed before the experiment to randomize the order of the categories. This randomization guarantees a fair distribution of categories and images among participants. The experimental platform follows a sequential and repetitive process as illustrated in Figure 1. (S1) The experimental platform presents the current category label. Participants can proceed to the next stage by pressing the space key. (S2) A fixation cross is shown at the center of the screen, ensuring attention is drawn when images are displayed. This fixation period lasts for 500 ms. (S3) The 50 images of this category are sequentially presented using the Rapid Serial Visual Presentation (RSVP) paradigm, which is commonly employed in psychological experiments. Each image is presented for a duration of 500 ms [12]. (S4) Random tests are conducted to verify the participant’s engagement in the experiment after the presentation. Data from categories for which participants fail the test will not be included in final analyses. The EEG signals of the participant will be captured and recorded continuously during the entire process. The program will cycle back to step S1 and display the next category, repeating this process until all the images have been presented.

### v2

# CrossPT-EEG: A Benchmark for Cross-Participant and Cross-Time Generalization of EEG-based Visual Decoding

@@ Abstract
Exploring brain activity in relation to visual perception provides insights into the biological representation of the world. While functional magnetic resonance imaging (fMRI) and magnetoencephalography (MEG) have enabled effective image classification and reconstruction, their high cost and bulk limit practical use. Electroencephalography (EEG), by contrast, offers low cost and excellent temporal resolution, but its potential has been limited by the scarcity of large, high-quality datasets and by block-design experiments that introduce temporal confounds. To fill this gap, we present CrossPT-EEG, a benchmark for cross-participant and cross-time generalization of visual decoding from EEG. We collected EEG data from 16 participants while they viewed 4,000 images sampled from ImageNet, with image stimuli annotated at multiple levels of granularity. Our design includes two stages separated in time to allow cross-time generalization and avoid block-design artifacts. We also introduce benchmarks tailored to non-block design classification, as well as pre-training experiments to assess cross-time and cross-participant generalization. These findings highlight the dataset’s potential to enhance EEG-based visual brain-computer interfaces, deepen our understanding of visual perception in biological systems, and suggest promising applications for improving machine vision models.
@@ 1 Introduction
Recent progress has been made in extracting clues from the human brain to inform advancements in AI and neuroscience, largely driven by the extensive use of functional magnetic resonance imaging (fMRI) (heeger2002does; logothetis2008we; logothetis2001neurophysiological) and magnetoencephalogram (MEG) (benchetrit2023brain) datasets. fMRI and MEG are widely used to investigate various cognitive functions, neurological disorders, and brain connectivity patterns (antonello2024scaling; ye2024relevance; toneva2022same; tang2024brain). Driven by the use of deep neural networks, particularly diffusion-based and transformer-based models, it is even possible to reconstruct human’s visual perceptions from fMRI or MEG recordings (takagi2023high; scotti2024reconstructing; ozcelik2023natural; cheng2023reconstructing). These successes are largely attributed to the availability of large-scale datasets, which offer comprehensive data essential to perform extensive studies and in-depth analyses (richards2019deep; lin2014microsoft). These models and large-scale datasets have opened new avenues for understanding the brain’s intricate functions and for developing advanced applications in brain-computer interfaces, neuroimaging, and beyond (st2023brain).
@@
In addition to fMRI and MEG, electroencephalography (EEG) is another vital tool in neuroscience research. EEG is easy to use, cost-efficient, and has high temporal resolution, making it a valuable tool for capturing rapid and real-time brain dynamics on the order of milliseconds (teplan2002fundamentals). EEG signals can be obtained non-invasively by placing electrodes on the scalp, making it a less intrusive method for monitoring brain activity. However, studies on visual perception with EEG signals are limited because of three challenges: (1) the lack of large-scale, high-quality EEG dataset; (2) existing EEG datasets typically featured coarse-grained image categories, lacking fine-grained categories; and (3) many existing datasets use block-design experiments, causing temporal effects that degrade data quality.
On the one hand, existing visual-EEG datasets such as spampinato2017deep are limited by a small number of participants and a restricted set of stimulus images. The limited data volume of current EEG datasets limits research findings’ statistical power and generalizability. For example, Things-EEG gifford2022large employs a Rapid Serial Visual Presentation (RSVP) paradigm with extremely short image presentation durations (e.g. 50 ms or 100 ms), where only early neural responses are captured, limiting the ability to study the full temporal dynamics of visual perception. On the other hand, the labels in existing EEG datasets are frequently coarse and lack the granularity needed for detailed analysis. Multi-granularity labels are essential because they allow for a more nuanced analysis at different levels of detail. For instance, labels can range from broad categories like “panda” or “golf ball” to more specific attributes like “Rottweiler” or “Samoyed”. Third, recent work shows a block-design experimental paradigm, in which EEG signals are affected by temporal effects both before and after stimuli, leading to biases in within-session classification results (li2020perils). These challenges underscore the necessity for new visual-EEG datasets and benchmarks that incorporate larger-scale, multi-granularity labels as well as cross-time experimental designs.
To address these challenges, we present CrossPT-EEG, a novel EEG benchmark specifically designed to promote research related to visual neuroscience, biomedical engineering, etc. As shown in Figure 1, CrossPT-EEG has a comprehensive dataset that includes EEG recordings from 16 subjects, with a total of 22 sessions, each exposed to 4,000 images sourced from the ImageNet-21k (ridnik2021imagenet). These images span 80 different categories, with 50 images per category. The dataset is structured to support multi-granularity analysis, with 40 categories dedicated to coarse-grained tasks and 40 to fine-grained. For 6 of the participants, we conducted Stage 2 experiment after a 7-day interval, adopting a non-block-design experimental paradigm. CrossPT-EEG incorporates classification experiments on non-block-design data, which avoids temporal confounds and thus offers a more reliable evaluation of EEG-based visual decoding. Furthermore, it integrates cross-time experiments that evaluate model generalization across sessions separated by several days, cross-participant experiments that assess robustness across individuals, and pre-training experiments that explore whether leveraging data from other participants can enhance cross-time recognition. Together, these tasks form a comprehensive benchmark designed to evaluate the robustness, generalization, and scalability of EEG-based visual decoding models, offering a solid foundation for both neuroscience and brain–computer interfaces.
We summarize the main contributions of this work as follows: 1) We propose CrossPT-EEG, the first EEG benchmark with multi-granularity semantic labels, designed for cross-participant and cross-time generalization of visual decoding from electroencephalogram. 2) We significantly scales up existing EEG-visual datasets in the number of subject sessions and total recording duration, providing high-quality and long-duration EEG segments for semantic-level visual understanding. 3) We establish benchmarks including classification tasks on non-block-design EEG data as well as pre-training experiments across time participants, thereby enabling systematic evaluation of model robustness and generalization.
@@ 2 Related Work
In this section, we review some datasets related to visual recognition and neuroscience and compare them with CrossPT-EEG, as shown in Table 1.
| Dataset | #Subjects | Modalities | Visual Stimuli | #Stimuli | Stimuli duration |
| GOD | |||||
| (horikawa2017generic) | 5 | fMRI | ImageNet | 1,200 | 1 s |
| NSD | |||||
| (allen2022massive) | 8 | fMRI | MS COCO | 10,000 | 3 s |
| DEAP | |||||
| (koelstra2011deap) | 32 | EEG, ECG | music video (1 min) | 40 | - |
| SEED | |||||
| (zheng2015investigating) | 15 | EEG | movie clips (4 min) | 15 | - |
| spampinato2017deep | 6 | EEG | ImageNet | 2000 | 500 ms |
| AMIGOS | |||||
| miranda2018amigos) | 40 | EEG, ECG | long/short videos | 20 | - |
| x2gf-5324-20 | 6 | EEG | ImageNet, Hollywood 2 | 2000+384 | 500 ms |
| Things EEG2 | |||||
| (gifford2022large) | 10 | EEG | Things | 16740 | 100 ms |
| Things EEG1 | |||||
| (grootswagers2022human) | 50 | EEG | Things | 22248 | 50 ms |
| EEG-SVRec | |||||
| (zhang2024eeg) | 30 | EEG, ECG | short videos | 2636 | - |
| EIT-1M | |||||
| (zheng2024eit) | 5 | EEG | CIFAR-10 | 60,000 | 50 ms |
| Alljoined | |||||
| (xu2024alljoined) | 8 | EEG | MS COCO | 10,000 | 300 ms |
| CrossPT-EEG (Ours) | 16+6 | EEG | ImageNet | 4,000 | 500 ms |
@@
Visual recognition is a cornerstone of computer vision, driven by datasets like ImageNet (ridnik2021imagenet), CIFAR (krizhevsky2009learning), and MS COCO (lin2014microsoft). Efforts to combine visual recognition with human have led to datasets like the SALICON dataset (jiang2015salicon), which extends MS COCO with eye-tracking data, enabling the study of visual attention and saliency through large-scale annotations. Neuroscience datasets utilizing fMRI have further enriched this field. The Generic Object Decoding dataset (horikawa2017generic) captures brain activity while subjects view and imagine objects, facilitating the decoding of mental images. The NSD (allen2022massive) is a large-scale fMRI dataset in visual neuroscience, recording high-resolution (1.8-mm) whole-brain 7T fMRI data from 8 subjects exposed to 9,000–10,000 color natural scenes from the MS COCO dataset over the course of a year.
@@
fMRI is renowned for its high spatial resolution, allowing researchers to obtain detailed images of brain activity by measuring changes in blood flow (logothetis2001neurophysiological). This capability makes fMRI particularly effective for identifying the specific brain regions involved in various cognitive and sensory tasks. In contrast, EEG offers several distinct advantages over fMRI. EEG is relatively easy to use and cost-efficient, with a straightforward setup that involves placing electrodes on the scalp to measure electrical activity. One of the most significant benefits of EEG is its exceptional temporal resolution, which captures neural dynamics on the order of milliseconds (teplan2002fundamentals). This makes EEG ideal for studying fast-occurring brain processes and real-time neural responses, providing insights into the timing and sequence of neural events (liu2021challenges). However, EEG signals collected using portable devices often have a low signal-to-noise ratio, which can complicate data analysis and reduce the accuracy of the results (kannathal2005characterization).
@@
Existing EEG datasets span a variety of research areas. The SEED (zheng2015investigating) focuses on emotion recognition with detailed EEG recordings from subjects exposed to various emotional stimuli. The BCI Competition IV datasets (zhang2012bci) provide EEG data for motor imagery tasks, while the TUH EEG Corpus (shah2018temple) is a large clinical EEG collection often used for benchmarking EEG data quality across different conditions. The DEAP (koelstra2011deap) collects EEG and peripheral physiological signals from 32 participants as they watch 40 one-minute music videos, providing comprehensive emotional responses. Similarly, the AMIGOS (miranda2018amigos) captures EEG and physiological responses from participants watching short video clips designed to evoke specific emotional states. In the realm of visual recognition, datasets like the EEG-Classification dataset (spampinato2017deep) involve 6 subjects viewing 2,000 images across 40 object classes from the ImageNet10k. The Things EEG (grootswagers2022human; gifford2022large) utilizes in the RSVP paradigm features extremely short image presentation durations (50/100 ms) and collects EEG-image pairs from large-scale datasets, making it a valuable resource for research in visual neuroscience. Another significant dataset is the EEG-SVRec (zhang2024eeg), which includes EEG recordings from 30 participants interacting with short videos, aiming to capture detailed affective experiences.
@@ 3 Dataset Construction
During the data collection process of our user study, participants are presented with a visual stimuli dataset containing 4000 natural images from ImageNet21k. Throughout this process, we continuously record their EEG signals. The whole experimental process is carried out in the laboratory environment. This section describes the entire process of CrossPT-EEG dataset construction.
@@ 3.1 Participants
We enlist a total of 16 participants via social media, including 10 males and 6 females. These participants are all college students aged between 21 and 27, with an average age of 24.06 and a standard deviation of 1.69. Their majors encompass computer science, mechanical engineering, chemistry, and environmental engineering, and they range from undergraduate to postgraduate levels. All participants are right-handed and assert their proficiency in utilizing image search engines in their daily routines. Each participant dedicates approximately 2 hours to complete the experiment each stage, which includes 30 minutes for equipment setup and task instructions. Before the experiment, participants are informed of a compensation of US$11.8 per hour upon completion, to ensure the quality of the data collected for the study.
@@ 3.2 Stimuli Dataset
Among all categories, the first half is consistent with the EEG-Classification dataset (spampinato2017deep), comprising 40 significantly distinct categories from ImageNet1k. We treat these as coarse-grained tasks. The latter 40 categories are designed as a fine-grained task, divided into 5 groups with 8 categories each. The categories within the same group share the same parent node in WordNet, and each category label is either a leaf node or a sub-leaf node in WordNet. This selection ensures that the chosen categories represent similar granularity while avoiding overly obscure categories, thereby minimizing potential biases in the experimental results. For instance, coarse-grained categories include items such as African elephants, pandas, mobile phones, golf balls, bananas, and pizzas. Under the parent node “musical instruments”, the fine-grained categories include accordions, cellos, flutes, oboes, snare drums, and trombones.
@@ 3.3 Procedure
Before engaging in the user study, participants are required to fill out an entry questionnaire and sign a consent about the protection of privacy security. They will receive an orientation regarding the primary tasks and operational procedures. Additionally, they will be notified of their right to withdraw from the study at any point. Before the main trials, participants will undergo a series of training trials designed to familiarize them with the procedures of formal experiments.
Before each experiment, every participant is required to select a unique random seed, which is used to randomize both the order of the categories and the order of images within each category. This randomization is applied in both experimental stages, ensuring a fair distribution of categories and images across participants. For the same participant, a different seed is used for Stage 1 and Stage 2. The experimental-platform code is publicly available in the code repository. The experimental platform follows a sequential and repetitive process as illustrated in Figure 1. (S1) The experimental platform presents the current category label. Participants can proceed to the next step by pressing the space key. (S2) A fixation cross is shown at the center of the screen, ensuring attention is drawn when images are displayed. This fixation period lasts for 500 ms. (S3) The images of this category are sequentially presented using the Rapid Serial Visual Presentation (RSVP) paradigm, which is commonly employed in psychological experiments. Each image is presented for a duration of 500 ms (kaneshiro2015representational), with a total number of . (S4) Random tests are conducted to verify the participant’s engagement in the experiment after the presentation. Specifically, at a random moment after completion of a category block, the participant is asked to answer which category was just presented. Data from categories for which participants fail the test will not be included in final analyses. The EEG signals of the participant will be captured and recorded continuously during the entire process. The program will cycle back to step S1 and display the next category, repeating this process until all the images have been presented.
Our experiment consists of two stages. The first stage follows the same setup as spampinato2017deep where images from each category are presented consecutively, with as shown in Figure 1. All of 16 participants took part in this stage. However, as li2020perils pointed out, the experimental results under the paradigm of such block-design may be influenced by temporal effects when using shuffled training and test sets. Therefore, we conducted a second stage of the experiment, with and random shuffling. Six participants participated in this phase. Stage 2 was conducted at least seven days after Stage 1. The 7-day interval was intended to prevent potential biases arising from memory effects (fisher2018patterns). Ultimately, the dataset we construct includes the EEG signals of participants exposed to each image visual stimulus in each valid session, along with the corresponding category’s wnid and the image’s index in ImageNet21k.